OpenStudy (msmr):

Solve by completing the square: 2x^2 - 8x + 3 = 0

OpenStudy (anonymous):

\[x=\sqrt{\sqrt{29/2}+4}\]

OpenStudy (mathmate):

2x^2 - 8x + 3 = 0 x^2-4x=-3/2 (x-2)^2=-3/2+4=5/2 x-2=+/- sqrt(5/2) x=2+/- sqrt(5/2)